OpenStudy (anonymous):

True or False: The factors of x(x + 8) - 4(x + 8) after using the grouping method is 2(x + 8)(x - 4).

OpenStudy (ajprincess):

can u factorise xy-4y? @AntoinetteWillman

OpenStudy (anonymous):

No.

OpenStudy (ajprincess):

hmm k. what is common to both xy and 4y?

OpenStudy (anonymous):

y?

OpenStudy (ajprincess):

ya good:) nw what remains when u take it out from xy+4y

OpenStudy (anonymous):

x+4

OpenStudy (ajprincess):

ya good:) so xy+4y=y(x+4)

OpenStudy (ajprincess):

Nw x(x + 8) - 4(x + 8)=?

OpenStudy (anonymous):

So its true right?

OpenStudy (ajprincess):

nope.

OpenStudy (anonymous):

Lol oops thank you for your help and time.

OpenStudy (ajprincess):

do u knw y it is nt true?

OpenStudy (anonymous):

Because they are equivalent to each other.

OpenStudy (ajprincess):

nope. x(x + 8) - 4(x + 8)=(x-4)(x+8) but the answer given is 2(x-4)(x+8)